40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10378     Entry time: Wed Aug 13 19:23:09 2014
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PSL, Aux laser mode hop check 

This afternoon Q helped me put in some temporary PDs for checking for any mode hopping behavior in our 3 main lasers. 

Q helped me install PDA55s on each of the lasers (I did the ends, he did the PSL) so that we could do the mode hop temperature check.  For the Yend, I took the leakage transmission through the first Y1 steering mirror after the laser. This beam was dumped, so I replaced the dump with a PDA55. For the Xend, the equivalent mirrors are too close to the edge of the table, so I put in a spare Y1, and reflect most of the light to a beam dump.  The leakage transmission then goes to a PDA55.  Note that for both of these cases, no alignment of main laser path mirrors was touched, so we should just be able to remove them when we're through.  For the PSL, I believe that Q took the rejected light from one of the PBSes before the PMC. 

The end temporary PDs are using the TRX / TRY cables, so we will be looking at the C1:LSC-TR[x,y] channels for the power of the end lasers.  The PSL's temporary PD is connected to the PMC REFL cable.  For the end PDs, since I had filter banks available, I shuttered the end lasers and removed the dark offset.  I then changed the gains to 1, so the values are in raw counts.  The usual transmission normalization gains are noted in one of the control room notebooks.

I did a slow ezcastep and ramped the temperature of all 3 lasers over about an hour.  Since we usually use the PSL around FSS slow slider value of zero, I swept that from -10 to +10.  Since we usually use the Xend laser at around 10,000 counts, I swept that from 0 to 20,000.  For the Yend laser, it is usually around -10,000 counts, so I swept it from -20,000 to 0.  ezcastep -s 0.2 C1:ALS-X_SLOW_SERVO2_OFFSET +1,20000 C1:ALS-Y_SLOW_SERVO2_OFFSET +1,20000 C1:PSL-FSS_SLOWDC +0.001,20000

I was looking for something kind of similar to what Koji saw when he did this kind of sweep for the old MOPA (elog #2008), but didn't see any power jumps that looked suspicious.

Here is the PSL:

ModeHopCheck_PSL_13Aug2014.pdf

The Xend:

ModeHopCheck_Xend_13Aug2014.pdf

And the Yend:

ModeHopCheck_Yend_13Aug2014.pdf

ELOG V3.1.3-